please help me in finding x, y and z in the linear equation
2x-5y=2
-4x+3z=5
3y+4z=12
thank you :)

Answer by Alan3354(69443)

please help me in finding x, y and z in the linear equation
2x-5y=2
-4x+3z=5
3y+4z=12
===========
Multiply the 1st eqn by 2, then add to the 2nd eqn.
That leaves 2 eqns in y & z, no x.